You are given the following information on Parrothead Enterprises:


Debt:
9,300 6.5 percent coupon bonds outstanding, with 22 years to maturity and a quoted price of 104.75. These bonds pay interest semiannually and have a par value of $1,000.


Common stock:
240,000 shares of common stock selling for $64.80 per share. The stock has a beta of .93 and will pay a dividend of $3.00 next year. The dividend is expected to grow by 5.3 percent per year indefinitely.


Preferred stock:
8,300 shares of 4.65 percent preferred stock selling at $94.30 per share. The par value is $100 per share.


Market:
11.7 percent expected return, risk-free rate of 3.75 percent, and a 23 percent tax rate.


Calculate the company's WACC. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1
Type of Capital Market Value MV weights Specific Cost of Capital Weighted Cost of Capital
Bonds $ 974,175 0.05628 0.04697 0.002643
Common Stock 15,552,000 0.89850 0.111435 0.100124
Preferred Stock 782,690 0.04522 0.049311 0.002230
$17,308,865 1.00000 0.104997

WACC : 10.50 %

After tax cost of debt = 6.1 % x ( 1 - 0.23 ) = 4.697 % or 0.04697

Cost of common stock = 3.75 + 0.93 ( 11.7 - 3.75 ) = 11.1435 % or 0.111435

Cost of preferred stock = 4.65 / 94.30 = 4.9311 % or 0.049311